When Ricky was 10 years old, his father was 45. How many years must pass for Ricky to be 5 years older than one-fifth as old as his father?
Answer by mananth(16946)

let the number of years be x
Ricky's age after x years = x+10
Father's age after 5 years = x+45
x+10= 1/5(x+45)+5
multiply equation by 5
5(x+10)=(x+45)+25
5x+50=x+45+25
5x-x=25-50+45
4x=20
x=5 years
m.ananth@hotmail.ca